4n+6=8 the question is find four consecutive odd integers whose sum is 8. and that was my set up. i don't know how to find the answer.
Answer by piglet162431(90)

Let n be the first number
n The first number
+
n+2 The second
+
n+4
+
n+6 The last.
Combine.
4n+12 = 8
4n=-4
n=-1
-1, 1,3,5,
Check.
-1 + 1 +3+5
=0+8
=8